Exercise 8-6 Double-declining-balance depreciation LO P1 


Ramirez Company installs a computerized manufacturing machine in its factory at the beginning of the year at a cost of $60,000. The machine's useful life is estimated at 10 years, or 450,000 units of product, with a $6,000 salvage value. During its second year, the machine produces 49,000 units of product. 


Determine the machine's second-year depreciation using the double-declining-balance method. 


                           Double-declining-balance Depreciation 

Choose Factors:xChoose Factor(%)=Annual Depreciation Expense


x
=Depreciation expense
First year's depreciation
x
=
Second year's depreciation
x
=
Answer #1

Answer Depreciation expenses in the first year using double decline balance method =$12000.

Depreciation expenses in the second year using double decline balance method= $9600.

Explanation- Double Declining balance depreciation is calculated using the following formula:

Depreciation = Depreciation Rate * Book Value of Asset

Depreciation rate is given by the following formula:

Depreciation Rate = Accelerator *Straight Line Rate

Straight-line Depreciation Rate = 1/10 = 0.10 = 10%
Declining Balance Rate = 2*10% = 20%

Depreciation for first year = $60000 *20% = $12000

Book value at the end of first year = $60000-$12000= $48000

Depreciation for second year = $48000 *20% = $9600